# New 511 keV line data provides strongest sub-GeV dark matter constraints

### v3

@@ Appendix A Height scale of the free electron density
As explained above, in order to consider that different regions of the Galaxy will have a different rate (probability) of positronium formation, that also depends on the distribution of the ambient (free) electrons, we consider a scaling of the keV line emission following the NE2001 model Cordes & Lazio (2003a, b) for the free electron density distribution. In particular, we consider that the disk of the Galaxy will always contain a much larger density of free electrons in the ISM gas than diffuse positrons (i.e. there will be always more electrons available than positrons to form positronium states, thus being “saturated”, and the probability of positronium formation will mainly depend on the diffuse positron distribution). However, this assumption must be broken well above and below the disk, since the density of free electrons falls very quickly as the galactic height, , is increased, as shown by Fig. 3. In particular, in this figure we depict the electron density distribution, at the center of the Galaxy, as a function of the Galactic height, normalized to the electron density at the GC position. This is the scale factor that we apply to our evaluations of the keV line. This scaling is particularly important for very low DM masses, when the positrons lose energy very fast in the disk (but not outside the disk, since the density of gas producing these losses is much lower) and the fraction of positrons annihilated in flight (directly proportional to the gas density where the positrons propagate) becomes relevant.
@@ Appendix B Relevant timescales
In this appendix, we report a comparison of the longitude (left panels of Fig. 5) and latitude (right panels of Fig. 5) profiles predicted for DM annihilation, assuming different standard Galactic DM distributions, for DM masses of and MeV. In particular, we show the predictions obtained for the NFW profile (cyan lines), a contracted NFW profile (c-NFW), like the one that provides the best-fit to the Galactic Center Excess (GCE) Ackermann et al. (2017), which has a slope of (green lines), a Moore profile Moore et al. (1999) , which has a slope of (blue lines), and a cuspy profile, as an example of a distribution that accounts for spikes formed around the central black hole Sgr A*, taken from Refs. Gondolo & Silk (1999); Lacroix (2018); Balaji et al. (2023) (black lines).
We defer the full study of the compatibility of the different DM distributions with the SPI observations to a future work. However, this comparison allows us to see how the profiles are sensitive to the different choice of the Galactic DM distribution. In particular, we see that profiles that are not very steep (with slope ) are not able to produce a good simultaneous fit to the longitudinal and latitudinal SPI profiles for DM mass exceeding an MeV. However, a steeper profile would be able to provide relatively satisfactory fits for MeV masses. However, we remark that a high slope, like that adopted in the Moore profile, is already in tension with some astrophysical observations. This is because it can predict too much DM in the inner regions of galaxies and galaxy clusters to comply with rotation curve observations Spekkens & Giovanelli (2005), -ray observations of galactic clusters Walker et al. (2012), gravitational lensing observations Foëx et al. (2012); Vegetti & Vogelsberger (2014) and N-body simulations of CDM halos Navarro et al. (2004). In general, this leads to the conclusion that DM cannot be the dominant source of positrons in the Galaxy, but, instead, it is expected that a combination of different sources is producing the peaked keV emission profiles that SPI observes.
@@ Appendix D Annihilation and decay profiles for different channels
In this appendix, we report the predicted latitude and longitude profiles of the keV emission from DM decay and annihilation. In particular, we show in Fig. 6 a comparison of profiles obtained assuming DM annihilation through the (top panels), (middle panelbos) and direct (bottom panels) channels for different DM masses (from twice the mass of the final state, around MeV, to GeV, in the case of the and channels, and from MeV to GeV for the direct channel). The left panels illustrate the predicted longitude profiles of the keV emission and the right panels the predicted latitude profiles. These comparisons allow us to see that the data points corresponding to the higher longitude and latitude angles (in absolute value) are the most constraining ones, which explains why the bounds derived from this dataset are much stronger than those derived from Ref. Vincent et al. (2012) (besides the fact that these measurements are reported after double the duration of event collection). As we see, the predicted longitude profiles show a clear trend: the lower the mass of the DM particle, the more concentrated the keV emission around the GC. This is expected, since a lower DM mass means a lower injection energy for the positrons, and the lower the energy of the positrons, the slower they diffuse. Hence lower energy positrons have a distribution more similar to the DM distribution (NFW in these calculations) and lose their energy closer to their injection point. As we pointed out in the previous section, the predictions from a NFW DM distribution do not lead to a keV emission longitude profile as peaky as that observed by the SPI data even for the low DM masses that are accessible in the direct channel. In turn, as we see from the predicted latitude profiles, masses of around MeV reproduce the observed SPI latitude profile quite well. We remark that this is due to the scaling that accounts for the decay of the density of free electrons away from the GC, applied for first time in this work. Without accounting for this effect, we observe that the latitude profiles become much more diffuse (i.e. flatter profiles) than what is observed. In addition, for the latitude profiles, we see that the same trend wherein lower DM mass leads to a profile more concentrated around the Galactic Plane is observed above some tens of MeV. However, once the energy loss (mainly Coulomb and ionization energy losses) becomes the dominant process during the positron propagation, the keV emission at the Galactic plane starts to decay quickly, given that these particles suffer from severe energy losses and annihilation in flight losses (since this effect is proportional to the gas density, it does not affect the positrons above and below the Galatic plane). We have investigated that the minimum energy employed in these simulations ( eV) is low enough that our results do not change appreciably for a minimum energy of eV.

### v4

@@ Appendix A Height scale of the free electron density
As explained above, in order to consider that different regions of the Galaxy will have a different rate (probability) of positronium formation, that also depends on the distribution of the ambient (free) electrons, we consider a scaling of the keV line emission following the NE2001 model [19, 20] for the free electron density distribution. In particular, we consider that the disk of the Galaxy will always contain a much larger density of free electrons in the ISM gas than diffuse positrons (i.e. there will be always more electrons available than positrons to form positronium states, thus being “saturated”, and the probability of positronium formation will mainly depend on the diffuse positron distribution). However, this assumption must be broken well above and below the disk, since the density of free electrons falls very quickly as the galactic height, , is increased, as shown by Fig. 3. In particular, in this figure we depict the electron density distribution, at the center of the Galaxy, as a function of the Galactic height, normalized to the electron density at the GC position. This is the scale factor that we apply to our evaluations of the keV line. This scaling is particularly important for very low DM masses, when the positrons lose energy very fast in the disk (but not outside the disk, since the density of gas producing these losses is much lower) and the fraction of positrons annihilated in flight (directly proportional to the gas density where the positrons propagate) becomes relevant.
@@ Appendix B Relevant timescales
In this appendix, we report a comparison of the longitude (left panels of Fig. 5) and latitude (right panels of Fig. 5) profiles predicted for DM annihilation, assuming different standard Galactic DM distributions, for DM masses of and MeV. In particular, we show the predictions obtained for the NFW profile (cyan lines), a contracted NFW profile (c-NFW), like the one that provides the best-fit to the Galactic Center Excess (GCE) [1], which has a slope of (green lines), a Moore profile [54] , which has a slope of (blue lines), and a cuspy profile, as an example of a distribution that accounts for spikes formed around the central black hole Sgr A*, taken from Refs. [38, 46, 5] (black lines).
We defer the full study of the compatibility of the different DM distributions with the SPI observations to a future work. However, this comparison allows us to see how the profiles are sensitive to the different choice of the Galactic DM distribution. In particular, we see that profiles that are not very steep (with slope ) are not able to produce a good simultaneous fit to the longitudinal and latitudinal SPI profiles for DM mass exceeding an MeV. However, a steeper profile would be able to provide relatively satisfactory fits for MeV masses. However, we remark that a high slope, like that adopted in the Moore profile, is already in tension with some astrophysical observations. This is because it can predict too much DM in the inner regions of galaxies and galaxy clusters to comply with rotation curve observations [67], -ray observations of galactic clusters [71], gravitational lensing observations [33, 69] and N-body simulations of CDM halos [56]. In general, this leads to the conclusion that DM cannot be the dominant source of positrons in the Galaxy, but, instead, it is expected that a combination of different sources is producing the peaked keV emission profiles that SPI observes.
@@ Appendix D Annihilation and decay profiles for different channels
In this appendix, we report the predicted latitude and longitude profiles of the keV emission from DM decay and annihilation. In particular, we show in Fig. 6 a comparison of profiles obtained assuming DM annihilation through the (top panels), (middle panelbos) and direct (bottom panels) channels for different DM masses (from twice the mass of the final state, around MeV, to GeV, in the case of the and channels, and from MeV to GeV for the direct channel). The left panels illustrate the predicted longitude profiles of the keV emission and the right panels the predicted latitude profiles. These comparisons allow us to see that the data points corresponding to the higher longitude and latitude angles (in absolute value) are the most constraining ones, which explains why the bounds derived from this dataset are much stronger than those derived from Ref. [70] (besides the fact that these measurements are reported after double the duration of event collection). As we see, the predicted longitude profiles show a clear trend: the lower the mass of the DM particle, the more concentrated the keV emission around the GC. This is expected, since a lower DM mass means a lower injection energy for the positrons, and the lower the energy of the positrons, the slower they diffuse. Hence lower energy positrons have a distribution more similar to the DM distribution (NFW in these calculations) and lose their energy closer to their injection point. As we pointed out in the previous section, the predictions from a NFW DM distribution do not lead to a keV emission longitude profile as peaky as that observed by the SPI data even for the low DM masses that are accessible in the direct channel. In turn, as we see from the predicted latitude profiles, masses of around MeV reproduce the observed SPI latitude profile quite well. We remark that this is due to the scaling that accounts for the decay of the density of free electrons away from the GC, applied for first time in this work. Without accounting for this effect, we observe that the latitude profiles become much more diffuse (i.e. flatter profiles) than what is observed. In addition, for the latitude profiles, we see that the same trend wherein lower DM mass leads to a profile more concentrated around the Galactic Plane is observed above some tens of MeV. However, once the energy loss (mainly Coulomb and ionization energy losses) becomes the dominant process during the positron propagation, the keV emission at the Galactic plane starts to decay quickly, given that these particles suffer from severe energy losses and annihilation in flight losses (since this effect is proportional to the gas density, it does not affect the positrons above and below the Galatic plane). We have investigated that the minimum energy employed in these simulations ( eV) is low enough that our results do not change appreciably for a minimum energy of eV.
@@
Erratum version
Abstract
We correct the results obtained using the diffuse -ray emission from Xmm-Newton, in light of new results from Ref. [4], that demonstrated that the dataset employed was misevaluated. We also update our calculation of the keV emission from evaporating PBHs, which leads to slightly more conservative constraints as well.
Main text:
Erratum: New 511 keV line data provides strongest sub-GeV dark matter constraints
Pedro De la Torre Luque
Shyam Balaji Joseph Silk
In the original paper [25], our calculation of the 511 keV flux inadvertently overcounted the number of positrons contributing to positronium formation. By integrating the positron flux over all energies, we included high-energy positrons that typically escape or annihilate in-flight before thermalizing, and therefore do not produce 511 keV photons in the region studied. As a result, 511 keV flux was over predicted, and the required cross sections/decay lifetimes were quoted too stringently. More precisely, our expression for the total diffuse positron flux ( in Eq. (3) of [25]) was defined as the integral of the positron flux over all positron energies. Instead, as shown in Ref.[45], a correct evaluation of the 511 keV flux requires taking the flux of positrons at the thermal energy of the interstellar gas where positronium forms, which, assuming warm gas, can be taken to be eV (see Ref. [45] for further discussion). The corrected calculation of the keV flux per solid angle is the following
where is the fraction of positronium decays corresponding to (singlet) para-positronium states) contributing to the keV line signal, the factor of accounts for the production of two photons in these decays, and is the energy-integrated emissivity of 511 keV photons, defined as:
where is the positron flux at the thermal energy, the positronium formation cross sections are the charge-exchange cross section with hydrogen [41]. The electron density, , is set to cm-3 in the Galactic plane and having the vertical dependence of free electron density as described in the original paper.
This change in our estimations makes the spatial profile of the 511 keV line follow almost exactly our previous evaluation, but it significantly varies the normalization (i.e. the annihilation or decay rates) of the predicted signals. In fact, we generally find larger deviations in the normalization of the signals for larger DM masses – a MeV DM particle requires a factor larger cross section than in our previous estimation (in case of decay, a factor lower decay rate), at MeV a factor ( for decay), at MeV a factor ( for decay), and at MeV a factor ( for decay). These factors are roughly the same for the different annihilation/decay channels. For brevity, we do not repeat every numerical value quoted in the original text, the updated constraints should be read with the revised normalization factors given above and in Figs. 1–2.
We show the updated version of Fig. 1 in the original paper [25] in Fig. S1. As it can be seen, the main difference with respect to the original version is in the annihilation rate required to match the observed fluxes.
We provide in Fig. S2 the updated DM bounds for annihilation (left panel) and decay (right panel) with this updated implementation. Note that these panels do not contain the XMM-Newton bounds, previously shown. This is because an updated analysis of the XMM-Newton diffuse data [4] resulted in constraints that are too weak, in comparison to the ones shown in Fig. S2.
The signal morphology is essentially unchanged, and the 511 keV line continues to provide highly competitive constraints on sub-GeV dark matter. The corrected implementation affects only the normalization of the limits: cross sections required for annihilation are larger and lifetimes are smaller than previously quoted. This systematically weakens the numerical bounds, but they remain among the strongest astrophysical constraints on light dark matter, particularly at low masses.